OpenStudy (anonymous):

how to solve arctan[tan(7pi /4)]

OpenStudy (mathmagician):

7pi/4

OpenStudy (mathmagician):

atan and tan are inverse functions

OpenStudy (anonymous):

but why tan (7pi/4) =-1

OpenStudy (mathmagician):

so you can simplify 7pi/4 and write that 7pi/4=-pi/4 or 3pi/4

OpenStudy (anonymous):

yes tan and arctan are inverse of each other

OpenStudy (anonymous):

so the answer will change ?!!

OpenStudy (mathmagician):

no, it will still be -pi/4

OpenStudy (anonymous):

find a place on the unit circle in the interval \[[-\frac{\pi}{2},\frac{\pi}{2}]\] where the tangent is also -1

OpenStudy (anonymous):

\[\tan(\frac{7\pi}{4})=\frac{\sin(\frac{7\pi}{4})}{\cos(\frac{7\pi}{4})}=-1\]